# Cosmological Simulations of Stellar Halos with Gaia Sausage–Enceladus Analogues: Two Sausages, One Bun?

### v1

@@ 6 Conclusion
Though we find it difficult to distinguish stars’ origins in 2D projections of kinematic spaces, machine learning techniques may be able to take advantage of the high dimensionality of the available parameter space to effectively classify stars according to their galactic progenitors. Such an algorithm would require a large amount of training data. This study is already a step towards a large statistical sample of GSE-like events, and as simulations continue to progress, sample sizes will continue to grow.
@@
Furthermore, though kinematic spaces such as $v_{r}-v_{\phi}$ and $E-L_{z}$ are ineffective at distinguishing merger events from one another, other parameters such as orbital frequencies may hold more discerning power (Gómez & Helmi, 2010). This requires detailed modeling of the MW’s gravitational potential, and efforts in this regard are well underway (e.g. Reino et al., 2021; Garavito-Camargo et al., 2021; Palau & Miralda-Escudé, 2023; Ibata et al., 2024). Further, the GSE itself influences the MW’s potential, which has observational consequences (e.g. Han et al., 2022, 2023a, 2023b). Including these efforts into the modeling of GSE debris will further aid in understanding its origin and composition.
@@
Crucially, star formation histories and chemical abundances do uncover systematic shifts between single-merger GSE debris and two-merger GSE debris, shown in §5. Single mergers are typically accreted later than the pairs, and correspondingly they have larger masses (both in terms of their stellar mass and their total mass), a longer duration to their star formation, and a slight metal enrichment relative to the pairs. These general trends do not, however, necessarily imply that these cases are observationally distinguishable, especially given that the largest shifts are seen in the distributions of stellar ages, which have observational uncertainties on the order of gigayears (e.g. Chaplin et al., 2014; Gallart et al., 2019; Grunblatt et al., 2021; Johnson et al., 2023).
@@
Determining the details of the GSE’s SFH is, therefore, a productive step in understanding its true nature. The literature has progressed in this regard, with many studies (e.g. Helmi et al., 2018; Gallart et al., 2019; Bonaca et al., 2020; Montalbán et al., 2021; Grunblatt et al., 2021; Feuillet et al., 2021; Hasselquist et al., 2021; Johnson et al., 2023; Horta et al., 2024; Ernandes et al., 2024) determining the ages of GSE stars and increasingly detailed accounts of GSE star formation. The general consensus of this body of work is that the peak of GSE star formation was $\sim 10$ Gyr ago, with a tail which continues for $\gtrsim 2$ Gyr as the dwarf fell into the MW and slowly stopped forming stars. The extended duration of star formation is consistent with a gradual infall with large pericenter—cf. Figure 7 and the simulations of Naidu et al. (2021)—or multiple smaller mergers—cf. Donlon et al. (2022); Rey et al. (2023); Orkney et al. (2023). There is still more work to be done in this direction: as a greater number of stars have their ages robustly determined, the time resolution of GSE star formation will be improved, giving deeper insight into the extragalactic source of this debris.
@@ Acknowledgments
The authors would like to acknowledge Andrea Caputo, Andreia Carrillo, Akaxia Cruz, James Johnson, Kassidy Kollmann, Matthew Orkney, Jonah Rose, and Risa Wechsler for helpful conversations. DF and ML are supported by the Department of Energy (DOE) under Award Number DE-SC0007968. ML is also supported by the Simons Investigator in Physics Award. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). LN is supported by the Sloan Fellowship, the NSF CAREER award 2337864, and NSF award 2307788. LH acknowledges support by the Simons Collaboration on “Learning the Universe.” The computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University. The IllustrisTNG simulations were undertaken with compute time awarded by the Gauss Centre for Supercomputing (GCS) under GCS Large-Scale Projects GCS-ILLU and GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS), as well as on the machines of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.
Appendix A The Milky Way Analogues
The snapshot-99 Subfind IDs of the 98 Milky Way (MW) analogues studied in this work are given below. The criteria by which these are selected are provided in §2.2.
@@ Appendix B Technical Details for Determining Accretion Histories
The following describes the modifications made to Sublink to obtain the merger tree for the Milky Way (MW) analogues (§B.1) and the details of how stars are associated with mergers (§B.2). Together, these provide implementation details for the methods described in §2.
B.1 Milky Way Merger Trees
The snapshot-by-snapshot history of the MW analogues can be largely determined via the Sublink algorithm (Rodriguez-Gomez et al., 2015) and corresponding “merger tree” catalogues provided by Nelson et al. (2019). At a given snapshot, Sublink identifies candidate “descendant” Subfind halos in the subsequent snapshot which share particles with this “progenitor” halo. The unique choice of descendant—which is, ideally, either the same physical structure as the progenitor or a larger halo into which the progenitor merges—is chosen from this list of candidates, with priority given to those halos that contain the most-bound particles from the progenitor.
From this process, a full tree is built up, with “branches” consisting of a halo, its descendants in subsequent snapshots, and its progenitors in previous snapshots. An example of such a tree is shown in Figure B1, in which the host halo (i.e., the MW analogue, henceforth simply called “the MW” for ease of discussion) has a branch determined by Sublink which is shown in gray, and the largest mergers are shown in various colors, with the opacity indicating distance from the MW; the markers are fully opaque when the merger is at or within $R_{\mathrm{dyn}}$. This MW experiences a merger $\sim 1.5$ Gyr ago with a halo (the darker-blue markers) that has mass comparable to that of the MW itself. Such a violent merger is difficult to resolve, both in terms of Subfind isolating the individual halos from each other and further in terms of Sublink connecting the halos from each snapshot robustly into a merger tree. In this case, the halos of the MW branch dip in mass by over 1 dex, while the merging object increases in mass by a similar amount. At an earlier point in the MW’s evolution, approximately 12 Gyr ago, another similarly large merger occurred. In this case, the MW halos identified by Subfind again experience a jump in $M_{\mathrm{dyn}}$ of approximately 1 dex.
@@ B.1 Milky Way Merger Trees
Figure B1: An example merger tree for one MW analogue with ID 394621. The horizontal axis is the time corresponding to each snapshot, and the vertical axis is the mass of the halos in the merger tree at each snapshot. The MW analogue’s branch from Sublink is shown in solid gray and the patched branch used in this work is in dashed black. The largest mergers are shown with colored markers, where the opacity corresponds to the distance from the host halo: the markers are opaque when the merger is within the MW analogue’s $R_{\mathrm{dyn}}$. The snapshots on the Sublink branch marked with a green checkmark are those which are verified as correct (see text), while the snapshots marked with a red “X” are those where the halo ID used in this work is changed from the ID provided by Sublink.
@@
Jumps in halo mass which occur this rapidly may be indicative of cases where many particles previously bound to the halos of the MW branch instead get attributed to the halos of the merger’s branch. To mitigate these cases, we modify the MW branch: if the total bound mass $M_{\mathrm{dyn}}$ has a significant ($>0.2$ dex) increase between two snapshots, this is taken as an indication that such a misidentification has taken place. We assume that the higher-mass halo on the MW branch, in the snapshot closer to $z=0$, is correctly identified, and then double-check that the lower-mass halo in the prior snapshot is its true progenitor. This is done by taking the central black hole in the descendant halo and finding it in the preceding snapshot. The halo that contains the black hole in this snapshot is then labeled the main progenitor and used in place of the halo identified by Sublink. This process then recurses, starting at the most recent spurious snapshots and working backwards through time until either there are no such jumps in the mass or there is no longer a black hole present.
@@
The result of this process is shown in Figure B1: at the right edge of the figure, between the third-to-last and second-to-last snapshots, the MW halo experiences a large jump in mass. A black hole is found in the second-to-last snapshot, and we confirm that it resides in the MW halo from the preceding snapshot. This is indicated by the green check mark and the agreement between the Sublink tree (in gray) and the amended MW branch (in black). This case highlights the difficulties in merger tree algorithms: it is unlikely that the MW halo should truly be considered to have lost this much mass, but faithfully tracking the physical structures with a halo-finding algorithm is complicated, especially in large-mass-ratio mergers, and linking these halos together into a merger tree branch can be challenging. After this jump, the algorithm then recurses and finds no further jumps until very early on in cosmic history, at $t\sim 12$ Gyr ago. At this jump, a black hole exists and is used to re-select the MW halo. In this case, our algorithm rejects the low-mass halo found by Sublink (indicated with the red “X”). Instead, we find the black hole in a different halo, one which Sublink considers to be a merger rather than the MW itself. We use this halo as the MW rather than the halo chosen by Sublink. This process repeats for a total of five snapshots, where in each case the Sublink halo does not contain the MW’s black hole, so it is not used as part of the MW branch. These alterations result in a smoother mass evolution of the MW, without the sharp jump seen in the Sublink definition branch.
@@ B.2 Associating Stars to Mergers
This process will not in general fully characterize the merger history: halos which survive to $z=0$ will never have the MW as a descendant halo and therefore may not be in the MW’s tree. Therefore, we extend the Sublink tree using the data collected by tracking the individual stars (§2.3). We first identify stars which are not born in the MW analogue but which are never bound to any merger halo identified by the above procedure. We take Sublink trees for the halos which host these stars and isolate the branches corresponding to them following the above procedure, ensuring that the branch is tracked until merging with a larger halo, i.e. until the descendant halo has a different main progenitor.
With the MW analogue’s Sublink tree and the additional merger information added from the stars themselves, the full merger history of the MW analogue is known. As mentioned in the main text, stars are assigned to the merger to which they were bound for the greatest number of snapshots. In ambiguous cases, where a star spent the same number of snapshots in multiple mergers, preference is given to the merger with the largest peak mass. There is ambiguity for $1.25_{-1.08}^{+5.76}\%$ of ex situ stars in a given MW, and typically ($83_{-29}^{+15}\%$ of the time) these stars are found in the branches for only one snapshot each; these edge cases of ambiguous merger assignment are generally not stars which spend long periods of time bound to a merging halo.
Recall that we allow halos previously assigned to merger branches of the tree to be reassigned to the MW’s branch, which can happen when the MW halo momentarily has far fewer particles than in the subsequent snapshot. This happens especially in cases of mergers with large mass ratios and small pericentric distances, when particle assignment to halos and halo assignment to merger branches is difficult to establish. When this occurs, a single Sublink branch may be considered two mergers: the original branch is thought to be fully merged at pericenter, as its descendant halo is the MW, and if the merging halo is re-identified after its pericenter, this would be considered a second merger. A similar effect occurs if Sublink fails to connect halos between snapshots which are physically the same object. We checked that none of the radially-anisotropic (RA) pairs identified in this work are affected by this issue; they are all genuinely-independent mergers.
@@ Appendix C Additional Figures
The main body of this work highlights a few examples of RA debris in terms of kinematics (cf. Figure 4, which shows the distribution of $v_{r}$ versus $v_{\phi}$ for a selection of RA debris, and Figure 5, which shows the distribution of $E$ versus $L_{z}$ for the same mergers) and in terms of star formation history (cf. Figure 6) and chemical abundances (cf. Figure 8). For completeness, the rest of these distributions are provided below; see the main text for full descriptions of the data contained in the figures.
Figure C1: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 419618, which hosts a single RA merger.
Figure C2: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 454172, which hosts a single RA merger.
Figure C3: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 459558, which hosts a single RA merger.
Figure C4: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 504559, which hosts a single RA merger.
Figure C5: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 505100, which hosts a single RA merger.
Figure C6: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 512425, which hosts a single RA merger.
Figure C7: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 515296, which hosts a single RA merger.
Figure C8: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 520885, which hosts a single RA merger.
Figure C9: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 525002, which hosts a single RA merger.
Figure C10: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 528322, which hosts a single RA merger.
Figure C11: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 530330, which hosts a single RA merger.
Figure C12: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 537488, which hosts a single RA merger.
Figure C13: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 537941, which hosts a single RA merger.
Figure C14: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 540082, which hosts a single RA merger.
Figure C15: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 540920, which hosts a single RA merger.
Figure C16: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 541218, which hosts a single RA merger.
Figure C17: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 543729, which hosts a single RA merger.
Figure C18: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 550149, which hosts a single RA merger.
Figure C19: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 552414, which hosts a single RA merger.
Figure C20: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 569251, which hosts a single RA merger.
Figure C21: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 571633, which hosts a single RA merger.
Figure C22: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 392277, which hosts a RA pair.
Figure C23: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 394621, which hosts a RA pair.
Figure C24: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 446665, which hosts a RA pair.
Figure C25: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 459557, which hosts a RA pair.
Figure C26: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 495451, which hosts a RA pair.
Figure C27: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 516760, which hosts a RA pair.
Figure C28: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 534628, which hosts a RA pair.
Figure C29: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 535050, which hosts a RA pair.
Figure C30: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 538905, which hosts a RA pair.
Figure C31: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 543114, which hosts a RA pair.
Figure C32: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for MW 552581, which hosts a RA pair.
Figure C33: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for the second choice of RA pair in MW 392277.
Figure C34: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for the second choice of RA pair in MW 495451.
Figure C35: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for the second choice of RA pair in MW 534628.
Figure C36: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for the second choice of RA pair in MW 535050.
Figure C37: Versions of Figure 4, Figure 5, Figure 6, and Figure 8 for the second choice of RA pair in MW 552581.

### v2

@@ VI Conclusion
Though we find it difficult to distinguish stars’ origins in 2D projections of kinematic spaces, machine learning techniques may be able to take advantage of the high dimensionality of the available parameter space to effectively classify stars according to their galactic progenitors. Such an algorithm would require a large amount of training data. This study is already a step toward a large statistical sample of GSE-like events, and as simulations continue to progress, sample sizes will continue to grow.
@@
Furthermore, though kinematic spaces such as $v_{r}-v_{\phi}$ and $E-L_{z}$ are ineffective at distinguishing merger events from one another, other parameters such as orbital frequencies may hold more discerning power (Gómez and Helmi, 2010). This requires detailed modeling of the MW’s gravitational potential, and efforts in this regard are well underway (e.g. Garavito-Camargo et al., 2021; Reino et al., 2021; Palau and Miralda-Escudé, 2023; Ibata et al., 2024). Further, the GSE itself influences the MW’s potential, which has observational consequences (e.g. Han et al., 2022, 2023a, 2023b). Including these efforts into the modeling of GSE debris will further aid in understanding its origin and composition.
@@
Crucially, SFHs and chemical abundances do uncover systematic shifts between single-merger GSE debris and two-merger GSE debris, shown in Section V. Single mergers are typically accreted later than the pairs, and correspondingly they have larger masses (both in terms of their stellar mass and their total mass), a longer duration to their star formation, and a slight metal enrichment relative to the pairs. These general trends do not, however, necessarily imply that these cases are observationally distinguishable, especially given that the largest shifts are seen in the distributions of stellar ages, which have observational uncertainties on the order of gigayears (e.g. Chaplin et al., 2014; Gallart et al., 2019; Grunblatt et al., 2021; Johnson et al., 2023).
@@
Determining the details of the GSE’s SFH is, therefore, a productive step in understanding its true nature. The literature has progressed in this regard, with many studies (e.g. Helmi et al., 2018; Gallart et al., 2019; Bonaca et al., 2020; Feuillet et al., 2021; Grunblatt et al., 2021; Hasselquist et al., 2021; Montalbán et al., 2021; Johnson et al., 2023; Ernandes et al., 2024; Horta et al., 2024) determining the ages of GSE stars and increasingly detailed accounts of GSE star formation. The general consensus of this body of work is that the peak of GSE star formation was $\sim 10$ Gyr ago, with a tail that continues for $\gtrsim 2$ Gyr as the dwarf fell into the MW and slowly stopped forming stars. The extended duration of star formation is consistent with a gradual infall with large pericenter—see Figure 7 and the simulations of Naidu et al. (2021)—or multiple smaller mergers—see Donlon et al. (2022); Orkney et al. (2023); Rey et al. (2023). There is still more work to be done in this direction: as a greater number of stars have their ages robustly determined, the time resolution of GSE star formation will be improved, giving deeper insight into the extragalactic source of this debris.
@@ Acknowledgments
The authors would like to acknowledge Andrea Caputo, Andreia Carrillo, Akaxia Cruz, James Johnson, Kassidy Kollmann, Matthew Orkney, Jonah Rose, and Risa Wechsler for helpful conversations. D.F. and M.L. are supported by the Department of Energy (DOE) under award number DE-SC0007968. M.L. is also supported by the Simons Investigator in Physics Award. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). L.N. is supported by the Sloan Fellowship, the NSF CAREER award 2337864, and NSF award 2307788. L.H. acknowledges support by the Simons Collaboration on “Learning the Universe.” The computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University. The IllustrisTNG simulations were undertaken with compute time awarded by the Gauss Centre for Supercomputing (GCS) under GCS Large-Scale Projects GCS-ILLU and GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS), as well as on the machines of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.
Appendix A The Milky Way Analogs
The snapshot-99 Subfind IDs of the 98 Milky Way (MW) analogs studied in this work are given below. The criteria by which these are selected are provided in Section II.2.
@@ Appendix B Technical Details for Determining Accretion Histories
The following describes the modifications made to Sublink to obtain the merger tree for the MW analogs (Section B.1) and the details of how stars are associated with mergers (Section B.2). Together, these provide implementation details for the methods described in Section II.
B.1 MW Merger Trees
@@
Figure B1: An example merger tree for one MW analog with ID 394621. The horizontal axis is the time corresponding to each snapshot, and the vertical axis is the mass of the halos in the merger tree at each snapshot. The MW analog’s branch from Sublink is shown in solid gray, and the patched branch used in this work is in dashed black. The largest mergers are shown with colored markers, where the opacity corresponds to the distance from the host halo: the markers are opaque when the merger is within the MW analog’s $R_{\mathrm{dyn}}$. The snapshots on the Sublink branch marked with a green check mark are those that are verified as correct (see text), while the snapshots marked with a red cross are those where the halo ID used in this work is changed from the ID provided by Sublink.
The snapshot-by-snapshot history of the MW analogs can be largely determined via the Sublink algorithm (Rodriguez-Gomez et al., 2015) and corresponding “merger tree” catalogs provided by Nelson et al. (2019b). At a given snapshot, Sublink identifies candidate “descendant” Subfind halos in the subsequent snapshot that share particles with this “progenitor” halo. The unique choice of descendant—which is, ideally, either the same physical structure as the progenitor or a larger halo into which the progenitor merges—is chosen from this list of candidates, with priority given to those halos that contain the most-bound particles from the progenitor.
From this process, a full tree is built up, with “branches” consisting of a halo, its descendants in subsequent snapshots, and its progenitors in previous snapshots. An example of such a tree is shown in Figure B1, in which the host halo (i.e., the MW analog, henceforth simply called “the MW” for ease of discussion) has a branch determined by Sublink that is shown in gray, and the largest mergers are shown in various colors, with the opacity indicating distance from the MW; the markers are fully opaque when the merger is at or within $R_{\mathrm{dyn}}$. This MW experienced a merger $\sim 1.5$ Gyr ago with a halo (the darker blue markers) that has a mass comparable to that of the MW itself. Such a violent merger is difficult to resolve, both in terms of Subfind isolating the individual halos from each other and further in terms of Sublink connecting the halos from each snapshot robustly into a merger tree. In this case, the halos of the MW branch dip in mass by over 1 dex, while the merging object increases in mass by a similar amount. At an earlier point in the MW’s evolution, approximately 12 Gyr ago, another similarly large merger occurred. In this case, the MW halos identified by Subfind again experience a jump in $M_{\mathrm{dyn}}$ of approximately 1 dex.
@@
Jumps in halo mass that occur this rapidly may be indicative of cases where many particles previously bound to the halos of the MW branch instead get attributed to the halos of the merger’s branch. To mitigate these cases, we modify the MW branch: if the total bound mass $M_{\mathrm{dyn}}$ has a significant (>0.2>0.2 dex) increase between two snapshots, this is taken as an indication that such a misidentification has taken place. We assume that the higher-mass halo on the MW branch, in the snapshot closer to $z=0$, is correctly identified and then double-check that the lower-mass halo in the prior snapshot is its true progenitor. This is done by taking the central black hole in the descendant halo and finding it in the preceding snapshot. The halo that contains the black hole in this snapshot is then labeled the main progenitor and used in place of the halo identified by Sublink. This process then recurses, starting at the most recent spurious snapshots and working backward through time until either there are no such jumps in the mass or there is no longer a black hole present.
@@
The result of this process is shown in Figure B1: at the right edge of the figure, between the third-to-last and second-to-last snapshots, the MW halo experiences a large jump in mass. A black hole is found in the second-to-last snapshot, and we confirm that it resides in the MW halo from the preceding snapshot. This is indicated by the green check mark and the agreement between the Sublink tree (in gray) and the amended MW branch (in black). This case highlights the difficulties in merger tree algorithms: it is unlikely that the MW halo should truly be considered to have lost this much mass, but faithfully tracking the physical structures with a halo-finding algorithm is complicated, especially in large-mass-ratio mergers, and linking these halos together into a merger tree branch can be challenging. After this jump, the algorithm then recurses and finds no further jumps until very early on in cosmic history, at $t\sim 12$ Gyr ago. At this jump, a black hole exists and is used to re-select the MW halo. In this case, our algorithm rejects the low-mass halo found by Sublink (indicated with the red cross). Instead, we find the black hole in a different halo, one that Sublink considers to be a merger rather than the MW itself. We use this halo as the MW rather than the halo chosen by Sublink. This process repeats for a total of five snapshots, where in each case the Sublink halo does not contain the MW’s black hole, so it is not used as part of the MW branch. These alterations result in a smoother mass evolution of the MW, without the sharp jump seen in the Sublink definition branch.
@@ B.2 Associating Stars to Mergers
This process will not in general fully characterize the merger history: halos that survive to $z=0$ will never have the MW as a descendant halo and therefore may not be in the MW’s tree. Therefore, we extend the Sublink tree using the data collected by tracking the individual stars (Section II.3). We first identify stars that are not born in the MW analog but that are never bound to any merger halo identified by the above procedure. We take Sublink trees for the halos that host these stars and isolate the branches corresponding to them following the above procedure, ensuring that the branch is tracked until merging with a larger halo, i.e. until the descendant halo has a different main progenitor.
With the MW analog’s Sublink tree and the additional merger information added from the stars themselves, the full merger history of the MW analog is known. As mentioned in the main text, stars are assigned to the merger to which they were bound for the greatest number of snapshots. In ambiguous cases, where a star spent the same number of snapshots in multiple mergers, preference is given to the merger with the largest peak mass. There is ambiguity for $1.25_{-1.08}^{+5.76}\%$ of ex situ stars in a given MW, and typically ($83_{-29}^{+15}\%$ of the time) these stars are found in the branches for only one snapshot each; these edge cases of ambiguous merger assignment are generally not stars that spend long periods of time bound to a merging halo.
Recall that we allow halos previously assigned to merger branches of the tree to be reassigned to the MW’s branch, which can happen when the MW halo momentarily has far fewer particles than in the subsequent snapshot. This happens especially in cases of mergers with large mass ratios and small pericentric distances, when particle assignment to halos and halo assignment to merger branches is difficult to establish. When this occurs, a single Sublink branch may be considered two mergers: the original branch is thought to be fully merged at pericenter, as its descendant halo is the MW, and if the merging halo is re-identified after its pericenter, this would be considered a second merger. A similar effect occurs if Sublink fails to connect halos between snapshots that are physically the same object. We checked that none of the RA pairs identified in this work are affected by this issue; they are all genuinely independent mergers.
@@ Appendix C Additional Figures
The main body of this work highlights a few examples of RA debris in terms of kinematics (see Figure 4, which shows the distribution of $v_{r}$ versus $v_{\phi}$ for a selection of RA debris, and Figure 5, which shows the distribution of $E$ versus $L_{z}$ for the same mergers) and in terms of star formation history (see Figure 6) and chemical abundances (see Figure 8). For completeness, the rest of these distributions are provided online. [^14] https://github.com/folsomde/Stellar_Halos_with_GSEs, or on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.14829722 (catalog doi:10.5281/zenodo.14829722).